OpenStudy (lisa100):

PLEASE HELP, big test tomorrow!! given: h(x)=185x-4x^2, find the vertex

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \(-\frac{185}{2\times(-8)}\) whatever that is

OpenStudy (anonymous):

second coordinate is what you get when you replace \(x\) by the answer above